# The dark matter content of Milky Way dwarf spheroidal galaxies: Draco, Sextans and Ursa Minor

### v1

@@ Abstract
journal: ApJ
@@ 1 Introduction
The standard $\Lambda$ cold dark matter ($\Lambda$CDM) cosmological model has successfully predicted the distribution of large-scale structures in our Universe (e.g. Yang et al., 2004; Cole et al., 2005; Eisenstein et al., 2005; Henriques et al., 2012; Dawson et al., 2013; Han et al., 2015; Wang et al., 2016; Springel et al., 2018; DESI Collaboration et al., 2024a). On small scales and within dark matter halos, there have been discrepancies between predictions of the standard cosmological model and observations (see Bullock & Boylan-Kolchin, 2017, for a review). One of the most hotly debated issues is the core-cusp problem (e.g. Flores & Primack, 1994; Moore, 1994; Burkert, 1995; de Blok et al., 2001; Gentile et al., 2004; Gilmore et al., 2007; de Blok, 2010; Oh et al., 2011). Cold dark matter simulations predict a universal density profile for dark matter halos with inner slopes close to 1 (cusp). By contrast, observations of low surface brightness galaxies, gas-rich dwarfs and dwarf spheroidal galaxies (dSphs) sometimes indicate inner density slopes close to 0 (core).
To understand the core-cusp problem, different scenarios have been proposed. Explanations beyond the $\Lambda$CDM theory turn to alternative dark matter models, such as self-interacting dark matter (e.g. Spergel & Steinhardt, 2000; Rocha et al., 2013; Oman et al., 2015; Foot & Vagnozzi, 2015; Kaplinghat et al., 2016; Jiang et al., 2023) and fuzzy dark matter (e.g. Hu et al., 2000; Hui et al., 2017). Within the CDM framework, stellar feedback is believed to be the most promising explanation, that repeated supernovae explosions move dark matter to more extended orbits and lower their densities in central regions (e.g. Navarro et al., 1996a; Read & Gilmore, 2005; Read et al., 2019; Freundlich et al., 2020a; Li et al., 2023; Boldrini, 2021). Numerical simulations have shown supporting evidences for this explanation (e.g. Mashchenko et al., 2008; Pontzen & Governato, 2012, 2014). Observationally, Read et al. (2019) reported an anti-correlation between the inner dark matter densities of dwarf galaxies and their stellar-to-halo mass ratios, which supports the stellar feedback scenario from observations (Di Cintio et al., 2014). Hayashi et al. (2020) summarized that the relation between the inner dark matter density slopes of classical dwarf galaxies and their stellar-to-halo mass ratios generally agrees with recent hydrodynamical simulations (e.g. Tollet et al., 2016; Freundlich et al., 2020b; Lazar et al., 2020).
@@
However, the constraints on the inner dark matter densities for dSphs are still limited by statistical errors (Zhu et al., 2016a) and systematic uncertainties behind dynamical modeling (Genina et al., 2018; Wang et al., 2022, 2023). For example, Read et al. (2019) and Hayashi et al. (2020) both measured the inner dark matter densities for eight classical dSphs, utilizing spherical and axisymmetric Jeans analysis respectively, but their measurements are not fully consistent with each other. In fact, cuspy dSphs can be biased to be cored due to violations of steady-state and/or spherical assumptions (Genina et al., 2018; Wang et al., 2022). Besides, binary orbital motions can inflate the overall velocity dispersion, whereas deflate the very central densities of dSphs in dynamical modeling (Wang et al., 2023), though the change is not significant for classical dSphs.
@@
The Milky Way dSphs are close enough to collect kinematical data of individual member stars for dynamical modeling. Recent observations have revealed the co-existence of multiple stellar populations in these systems (e.g. Tolstoy et al., 2004; Battaglia et al., 2008; Amorisco & Evans, 2012a; Pace et al., 2020; Arroyo-Polonio et al., 2024). Many studies model the stellar kinematics of different populations simultaneously to constrain the inner dark matter distributions in several dSphs (Battaglia et al., 2008; Walker & Peñarrubia, 2011; Amorisco & Evans, 2012b; Agnello & Evans, 2012; Amorisco et al., 2013; Zhu et al., 2016a; Strigari et al., 2017; Hayashi et al., 2018; Pace et al., 2020). For example, Walker & Peñarrubia (2011) estimate the slopes of mass profiles for Fornax and Sculptor defined by masses enclosed within half-number radii of two stellar populations, and found that their results are consistent with cored dark matter halos. In this study, we use the chemodynamical model developed in Zhu et al. (2016a) and Zhu et al. (2016b), which extends the discrete axisymmetric Jeans Anisotropic Multi-Gaussian Expansion (jam) model (Watkins et al., 2013). Specifically, we will adopt both the single-population jam model and a chemodynamical model in our analysis.
@@
We mainly use spectroscopic data for member stars in dSphs from the Milky Way Survey (MWS) of the Dark Energy Spectroscopic Instrument (DESI). DESI is currently one of the leading multi-object spectrographs for wide-field surveys (Levi et al., 2013; DESI Collaboration et al., 2016a, b, 2022; Silber et al., 2023; Schlafly et al., 2023; DESI Collaboration et al., 2024b). The Early Data Release (EDR; DESI Collaboration et al., 2024c) and Data Release 1 (DR1) of the first year observation (DESI Collaboration et al., 2025a) have already been publically available. DESI primarily focuses on observations of distant galaxies for extra-galactic science, such as exploring the nature of dark energy and mapping the large-scale structure of our Universe (DESI Collaboration et al., 2024a; Adame et al., 2025a; DESI Collaboration et al., 2024d, e; Adame et al., 2025b; DESI Collaboration et al., 2024f, 2025b). DESI MWS also observes millions of stars within the Milky Way to better understand our Galaxy (Cooper et al., 2023). The Year-3 data of DESI MWS includes member stars in Draco, Sextans and Ursa Minor.
@@
The paper is organized as follows. We first introduce spectroscopic data from DESI MWS for three dSphs and member star selection in Section 2. The single-population and chemodynamical model are described in Section 3. We present our results in Section 4, including measurements of the Astrophysical factors for dark matter detections. We compare with previous measurements in Section 5, including discussions on the core-cusp problem and selection effects. We draw conclusions in the end (Section 6).
@@
Figure 1: Top: A comparison between the Draco member star LOSV measurements by DESI MWS ($y$-axis) and in Walker et al. (2023) ($x$-axis). The red line is the best fit with slope of 1, and its $y$-intercept is the offset between two observations (see the legend). Bottom: The histogram of the differences between the LOSV measurements by DESI MWS and Walker et al. (2023). The red dashed vertical line represents the offset between two observations.
@@ 2.1 The DESI Milky Way Survey
The Dark Energy Spectroscopic Instrument (DESI) is currently one of the leading multi-object spectrographs for wide-field surveys (DESI Collaboration et al., 2016a, b, 2022, 2024c). The survey is carried on the Mayall 4 meter telescope at the Kitt Peak National Observatory (Miller et al., 2024; Poppett et al., 2024), with 5,000 fibers allocated over a $3.2^{\circ}$ diameter field of view at the prime focus. Although primarily focusing on observations of distant galaxies, DESI also collects spectroscopic data for stars within our Milky Way. The DESI Milky Way Survey (MWS) mainly operates during bright time, which refers to nights with more significant moon light contamination (Cooper et al., 2023; Koposov et al., 2024). The main survey of MWS will observe approximately seven million stars. DESI MWS also has extra secondary targets, such as faint blue horizontal branch stars (BHBs) in addition to those observed in the main survey. Beyond main survey and secondary targets, DESI MWS has tertiary programs, with dedicated fibers to observe special targets such as individual dSphs and stellar streams. In this work, we utilize the internal Year-3 data of DESI MWS, which covers data from the first three-year of observations. Note the DESI Year-1 data, commissioning and Science Validation (SV) data (DESI Collaboration et al., 2024b), which are already published in DESI DR1 (Koposov et al., 2025) and EDR (DESI Collaboration et al., 2024c; Koposov et al., 2024), are included in the latest internal version of Year-3 data.
@@ 2.2 DESI observations of Draco, Sextans and Ursa Minor
[…]
@@
We combine the DESI MWS data for three dSphs with the spectroscopic data from Walker et al. (2023), which are based on Magellan/M2FS and MMT/Hectochelle observations. Since some stars have multi-epoch observations in this data set by Walker et al. (2023), for each star we adopt the mean value of the LOSV (vlos_mean) and metallicity (feh_mean) measurements. The LOSV measurements by DESI and by Walker et al. (2023), however, may subject to different zero points, and we use stars observed by both DESI and Walker et al. (2023) to correct for this systematic offset. We show the LOSVs of stars observed by both DESI and Walker et al. (2023) and in the footprint of Draco as an example in Figure 1. There is an overall good agreement between the LOSV measurements by DESI and by Walker et al. (2023), with the data points in the top panel go well through the red diagonal line, but the LOSV measured by Walker et al. (2023) of Draco is on average lower than that of DESI by about 1 km/s , which is also obvious in the histogram of the differences in the bottom panel. This offset is added to the measurements of Walker et al. (2023).
For these stars that exist in both DESI and Walker et al. (2023), we use the LOSV measured with the smaller error. In Figure 2, we show the LOSV error distributions of stars in Draco, observed by DESI or in Walker et al. (2023). In general, LOSV measurements from DESI have larger errors than those from Walker et al. (2023), but both of them are smaller than the intrinsic LOSV dispersion of Draco (black dashed vertical line, from Pace et al. (2022)), indicating that they can be used for proper dynamical modeling. For metallicity measurements, we also adopt the one with the smaller measurement error for stars existing in both DESI and Walker et al. (2023), and we have corrected the mean offset of metallicity measurements by Walker et al. (2023) from those of DESI.
@@ 2.4 Member star selection
To select actual member stars for the three dSphs, we utilize the membership catalog for dwarf galaxies by Pace et al. (2022). The membership probabilities for stars associated to known MW dwarf galaxies are calculated by Pace et al. (2022) according to their Gaia parallaxes and spatial distributions, with a projected Plummer model adopted as the prior for the spatial distribution. Specifically, we first select stars with the probability mem_fixed or mem_fixed_complete greater than 0.9 from the membership catalog of Pace et al. (2022). We cross match the sample of stars observed by DESI in the footprints of the three dSphs or in Walker et al. (2023) to these high probability member stars, to obtain our sample of spectroscopically observed member stars. The matching is based on a searching radius of 0.3 arcsec. In the end, we further include $3\sigma$ clippings to the LOSVs and metallicities to further discard stars that are at the tail of the LOSV and metallicity distributions, which are possible contaminations. For LOSV, we use the systemic LOSV and its dispersion in Table 1 to set the $3\sigma$ boundaries for each dSphs. For metallicity we use the mean and standard deviation of our samples to set the $3\sigma$ boundaries. We also discard stars beyond $2.5$ half-light radius from the center of each dSphs to further exclude possible contaminations and dwarf halo stars (e.g. Deason et al., 2022).
@@
In our analysis, the central coordinates and systemic velocities of each dwarf galaxy have been subtracted from the observed coordinates and velocities for each member star in Cartesian coordinates, to correct the perspective rotation (Feast et al., 1961). The central coordinates and systemic proper motions for three dSphs are provided in Table 1. The systemic LOSV of each dSph is the mean LOSV of its member stars selected above.
@@
For each individual member star, not only their LOSVs are used for dynamical model, but also their Gaia proper motions are used. However, at the distances of the three dSphs used in our analysis, the uncertainties in Gaia proper motion measurements are quite large. Nevertheless, we still incorporated proper motions in our dynamical modeling, with the observational uncertainties for both LOSV and proper motions properly incorporated in the likelihood function (see Section 3.1 below), which will be taken into account by the model. Notably, we report the total velocity dispersion when comparing the actual and model predicted velocity dispersions instead of the intrinsic velocity dispersions deconvolved from observational errors.
@@
Though the pulsating velocities of DESI RR Lyraes have been corrected (Medina et al., 2025a, b) [^2] The DESI MWS RRL targets are based on the Gaia Data Release 2 (DR2; Clementini et al., 2019) and the Pan-STARRS1 (PS-1; Sesar et al., 2017) catalogs., we want to eliminate the possibility that some faint RR Lyraes may still remain unidentified to inflate the velocity dispersions of our dSphs. For safety, we exclude all member stars along the horizontal branch from our sample to remove potential faint RR Lyraes. We do not exclude possible binary stars since the effect of binary stars to inflate the velocity dispersion of classical dSph is very minor (e.g. Hargreaves et al., 1996; Minor et al., 2010; Spencer et al., 2017; Wang et al., 2023). We provide in Table 2 the final number of member stars in the three dSphs after our selections above, with the number of stars observed by DESI or in Walker et al. (2023) provided as well.
dSph $\alpha_{0}$ $\delta_{0}$ $r_{\mathrm{h}}$ $\mu_{\alpha\star}$ $\mu_{\delta}$ $d$ $v_{\mathrm{los}}$ $\sigma_{\mathrm{los}}$
$\left[\mathrm{deg}\right]$ $\left[\mathrm{deg}\right]$ $\left[\mathrm{arcmin}\right]$ $\left[\mathrm{mas\ yr^{-1}}\right]$ $\left[\mathrm{mas\ yr^{-1}}\right]$ $\left[\mathrm{kpc}\right]$ $\left[\mathrm{km\ s^{-1}}\right]$ $\left[\mathrm{km\ s^{-1}}\right]$
Draco $260.0684$ $57.9185$ $9.67\pm 0.09$ $0.044$ $-0.188$ $76$ $-290.7$ $9.1\pm 1.2$
Sextans $153.2628$ $-1.6133$ $16.50\pm 0.10$ $-0.409$ $0.037$ $86$ $224.3$ $7.9\pm 1.3$
UMi $227.242$ $67.2221$ $18.30\pm 0.11$ $-0.120$ $0.071$ $76$ $-247.0$ $8.6\pm 0.3$
Table 1: A summary of the central sky coordinates in Equatorial coordinates ($\alpha_{0}$ and $\delta_{0}$), half-light radii ($r_{h}$), systemic proper motions ($\mu_{\alpha}\ast$ and $\mu_{\delta}$), distances ($d$), systemic LOSVs ($v_{\mathrm{los}}$), and intrinsic LOSV dispersions ($\sigma_{\mathrm{los}}$) for the three dSphs (Draco, Sextans and UMi) used in the analysis of this paper. All of the information are taken from Pace et al. (2022), except for the distances, which are taken from Muñoz et al. (2018).
Object $N_{\mathrm{tot}}$ $N_{\mathrm{DESI}}$ $N_{\mathrm{W23}}$
Draco 407 96 374
Sextans 440 240 368
UMi 1046 1003 416
@@
Table 2: Numbers of member stars in the three dSphs used in our analysis. $N_{\mathrm{tot}}$ is the total number of member stars after selections in Section 2.4. $N_{\mathrm{DESI}}$ and $N_{\mathrm{W23}}$ are the numbers of member stars which are observed by DESI and from Walker et al. (2023), respectively. There are overlaps between the observed stars by DESI and in Walker et al. (2023), and for those stars existing in both data set, we use the LOSV and metallicity measurements with smaller errors (see Section 2.3).

### v2

@@ Abstract
journal: ApJthanks: haoyang22@sjtu.edu.cnthanks: Corresponding author: wenting.wang@sjtu.edu.cn
@@ 1 Introduction
[…]
@@
However, the constraints on the inner dark matter densities for dSphs are still limited by statistical errors (Zhu et al., 2016a) and systematic uncertainties behind dynamical modeling (Genina et al., 2018; Wang et al., 2022, 2023). For example, Read et al. (2019), Kaplinghat et al. (2019), Hayashi et al. (2020) and Andrade et al. (2024) all measured the inner dark matter densities for eight classical dSphs in our Milky Way (MW), utilizing spherical/axisymmetric Jeans analysis, or the phase-space distribution function method, but part of their measurements are not fully consistent with each other. In fact, cuspy dSphs can be biased to be cored due to violations of model assumptions (Genina et al., 2018; Wang et al., 2022). Besides, binary orbital motions can inflate the overall velocity dispersion, whereas deflate the very central densities of dSphs in dynamical modeling (e.g. Pianta et al., 2022), though the change is not expected to be significant for classical dSphs (Wang et al., 2023).
@@
The MW dSphs are close enough to have their individual member stars observed for dynamical modeling (Battaglia et al., 2013; Battaglia & Nipoti, 2022). Recent observations revealed the co-existence of multiple stellar populations in these systems (e.g. Tolstoy et al., 2004; Battaglia et al., 2008; Amorisco & Evans, 2012a; Pace et al., 2020; Arroyo-Polonio et al., 2024). Many studies model the stellar kinematics of different populations simultaneously to constrain the inner dark matter distributions in several dSphs (Battaglia et al., 2008; Walker & Peñarrubia, 2011; Amorisco & Evans, 2012b; Agnello & Evans, 2012; Amorisco et al., 2013; Zhu et al., 2016a; Strigari et al., 2017; Hayashi et al., 2018; Pace et al., 2020). For example, Walker & Peñarrubia (2011) estimate the slopes of mass profiles for Fornax and Sculptor defined by masses enclosed within half-number radii of two stellar populations, and found that their results are consistent with cored dark matter halos.
In this study, we use the chemodynamical model developed in Zhu et al. (2016a) and Zhu et al. (2016b), which extends the discrete axisymmetric Jeans Anisotropic Multi-Gaussian Expansion (jam) model (Watkins et al., 2013). Specifically, we adopt both the single-population jam model and a chemodynamical model in our analysis. Because the total mass within the half-number radius [^1] The radius within which it includes half of the stars in the sample. has been proven to be a sweet point (e.g. Wolf et al., 2010; Walker & Peñarrubia, 2011; Wang et al., 2015; González-Samaniego et al., 2017; Wang et al., 2022), which is not sensitive to mild deviations from model assumptions, joint modeling of multiple populations has the advantage of providing multiply measured total masses within the half-number radii of different populations, hence may bring better constraints than the single-population model.
@@
We mainly use spectroscopic data for member stars in dSphs from the MW Survey (MWS) of the Dark Energy Spectroscopic Instrument (DESI). DESI is currently one of the leading multi-object spectrographs for wide-field surveys (Levi et al., 2013; DESI Collaboration et al., 2016a, b, 2022; Silber et al., 2023; Schlafly et al., 2023; DESI Collaboration et al., 2024b). The Early Data Release (EDR; DESI Collaboration et al., 2024c) and Data Release 1 (DR1) of its first year data (DESI Collaboration et al., 2025a) are publicly available. DESI primarily focuses on distant galaxies to explore the nature of dark energy and mapping the large-scale structure of our Universe (DESI Collaboration et al., 2024a; Adame et al., 2025a; DESI Collaboration et al., 2024d, e; Adame et al., 2025b; DESI Collaboration et al., 2024f, 2025b). DESI MWS also observes millions of stars within the MW and MW dSphs (Cooper et al., 2023). The Year-3 data of DESI MWS includes member stars in Draco, Sextans and Ursa Minor.
@@
This paper is organized as follows. We first introduce spectroscopic data from DESI MWS for three dSphs and member star selection in Section 2. The single-population and chemodynamical model are described in Section 3. We present our results in Section 4, including measurements of the Astrophysical factors for dark matter detections. We compare with previous measurements in Section 5, including discussions on the core-cusp problem and selection effects. Our conclusions are presented in Section 6.
@@
Figure 1: Top: A comparison between the Draco member star LOSV measurements by DESI MWS ($y$-axis) and Walker et al. (2023) ($x$-axis). The red line is the best fit with slope of 1, and its $y$-intercept is the offset between two observations (see the legend). Bottom: The histogram of the differences between the Draco member star LOSV measurements by DESI MWS and Walker et al. (2023). The red dashed vertical line represents the offset between two observations determined from the fit in the top panel.
@@ 2.1 The DESI Milky Way Survey
The Dark Energy Spectroscopic Instrument (DESI) is currently one of the leading multi-object spectrographs for wide-field surveys (DESI Collaboration et al., 2016a, b, 2022, 2024c). The survey is carried on the Mayall 4 meter telescope at the Kitt Peak National Observatory (Miller et al., 2024; Poppett et al., 2024), with 5,000 fibers allocated over a $3.2^{\circ}$ diameter field of view at the prime focus. Although primarily focusing on observations of distant galaxies, DESI also collects spectroscopic data for stars. The DESI MWS mainly operates during bright time, which refers to nights with more significant moon light contamination (Cooper et al., 2023; Koposov et al., 2024). The main survey of MWS has observed nearly 12 million stars in the first three years. DESI MWS also has extra secondary targets, such as faint blue horizontal branch stars (BHBs). Beyond main survey and secondary targets, DESI MWS has tertiary programs, with dedicated fibers to observe special targets such as individual dSphs and stellar streams. In this work, we utilize the internal Year-3 data of DESI MWS.
@@ 2.2 DESI observations of Draco, Sextans and Ursa Minor
So far, DESI MWS has observed stars in footprints of Draco, Sextans and Ursa Minor (UMi). The properties of them are summarized in Table 1. Nearly $85\%$ and $30\%$ of the data for Draco and Sextans are taken at Survey Validation (SV), with the remaining from the main survey after SV. The data for UMi are from the tertiary program. We use the line-of-sight velocity (LOSV) measurements in the heliocentric frame, and also metallicity measurements of observed stars output by the DESI MWS RVS pipeline. The readers can find more details in Cooper et al. (2023), Koposov et al. (2024) and Koposov et al. (2025). We select stars with $\texttt{RVS\_WARN}=0$ to ensure the robustness of the stellar model fit by the pipeline [^2] RVS_WARN is a bitmask flag output by the RVS pipeline, which indicates potential issues or warnings in the LOSV measurements and stellar parameter determinations. The first bit of RVS_WARN is set to 1 if the difference in the $\chi^{2}$ values between the best-fit stellar model and the continuum-only model fit is small. The second bit is set to 1 if the LOSV is within $5\,\mathrm{km/s}$ of the predefined velocity boundary (1500-1500 to $1500\,\mathrm{km/s}$). The third bit is set to 1, if the LOSV uncertainty exceeds $100\,\mathrm{km/s}$. A spectrum without any of these concerns is indicated by $\texttt{RVS\_WARN}=0$. For stars with repeated observations in different programs, we use the observation with the smallest error of the LOSV measurement. Note for the LOSV errors reported by RVS, there is an additional systematic component or error floor of $0.9\,\mathrm{km/s}$, as mentioned in Cooper et al. (2023). We add this systematic error in quadrature to the LOSV errors in our analysis.
@@
We combine the DESI MWS data for three dSphs with the spectroscopic data from Walker et al. (2023), which are based on Magellan/M2FS and MMT/Hectochelle observations. Since some stars have multi-epoch observations by Walker et al. (2023), for each star we adopt the mean value of the LOSV (vlos_mean) and metallicity (feh_mean) measurements. The LOSV measurements by DESI and by Walker et al. (2023), however, may subject to different zero points, and we use stars observed by both DESI and Walker et al. (2023) to correct for this systematic offset. We show the LOSVs of stars observed by both DESI and Walker et al. (2023) and in the footprint of Draco as an example in Figure 1. There is a good agreement between two sets of LOSVs Walker et al. (2023). But the LOSV measured by Walker et al. (2023) of Draco is on average lower than that of DESI by about 1 km/s (see the histogram in the bottom panel). For each dSph, we determine the offset between two measurements as in Figure 1, and add this offset to the corresponding measurements of Walker et al. (2023).
For stars existing in both DESI and Walker et al. (2023), we use the LOSV measured with the smaller error. In Figure 2, we show the LOSV error distributions of stars in Draco, observed by DESI or in Walker et al. (2023). In general, LOSV measurements from DESI have larger errors than Walker et al. (2023), but both of them are smaller than the intrinsic LOSV dispersion of Draco (black dashed vertical line, from Pace et al. (2022)). For metallicity measurements, we also adopt the one with the smaller measurement error for stars existing in both DESI and Walker et al. (2023), and we have corrected the offset of metallicity measurements by Walker et al. (2023) from those of DESI for each dSph.
@@ 2.4 Member star selection
The membership probabilities for stars associated to known MW dwarf galaxies are calculated by Pace et al. (2022) according to their Gaia parallaxes and spatial distributions, with a projected Plummer model adopted as the prior for the spatial distribution. Specifically, we first select stars with the probability mem_fixed or mem_fixed_complete greater than 0.9 from the membership catalog of Pace et al. (2022). We cross match the sample of stars observed by DESI in the footprints of the three dSphs or in Walker et al. (2023) to these high probability member stars, to obtain our sample of spectroscopically observed member stars, with a searching radius of 0.3 arcsec. In the end, we further include $3\sigma$ clippings to the LOSVs and metallicities to remove stars at the tail of the distributions. For LOSV, we use the systemic LOSV and its dispersion in Table 1 to set the $3\sigma$ boundaries. For metallicity we use the mean and standard deviation of our samples. We also discard stars beyond $2.5$ half-light radius from the center of each dSphs to exclude possible interlopers and dwarf halo stars (e.g. Deason et al., 2022).
@@
The central coordinates and systemic velocities of each dSph have been subtracted from the observed coordinates and velocities for each member star in Cartesian coordinates, to correct the perspective rotation (Feast et al., 1961). The central coordinates and systemic proper motions are provided in Table 1. The systemic LOSV is the mean LOSV of its member stars selected above.
@@
For each member star, we use their LOSVs and Gaia proper motions for dynamical modeling. However, at the distances of the three dSphs used in our analysis, the uncertainties in Gaia proper motion measurements are quite large. Nevertheless, we still incorporated proper motions in our dynamical modeling, with the observational uncertainties for both LOSVs and proper motions properly incorporated in the likelihood function (see Section 3.1 below). Notably, when comparing the observed and model-predicted velocity dispersions, we report the total velocity dispersions which include observational errors, instead of the intrinsic velocity dispersions deconvolved from them.
@@
Though the pulsating velocities of DESI RR Lyrae stars have been corrected (Medina et al., 2025a, b) [^3] The DESI MWS RRL targets are based on the Gaia Data Release 2 (DR2; Clementini et al., 2019) and the Pan-STARRS1 (PS-1; Sesar et al., 2017) catalogs., we want to eliminate the possibility that some RR Lyrae stars may still remain unidentified to inflate the velocity dispersions of our dSphs. We thus exclude all stars along the horizontal branch. We do not exclude possible binary stars since the effect of binary stars to inflate the velocity dispersion of classical dSph is minor (e.g. Hargreaves et al., 1996; Minor et al., 2010; Spencer et al., 2017; Wang et al., 2023). Besides, member stars in UMi has multi-epoch observations in DESI (Qiu, et al., in preparation), based on which we have tried to exclude stars with large LOSV variabilities, but the change in LOSV dispersion is negligible. We provide in Table 1 the final number of member stars in the three dSphs after our selections above, with the number of stars observed by DESI or in Walker et al. (2023) provided as well.
Property Draco Sextans UMi
$\alpha_{0}$ [deg] 260.0684 153.2628 227.242
$\delta_{0}$ [deg] 57.9185 1.6133-1.6133 67.2221
$r_{\mathrm{h}}$ [arcmin] $9.67\pm 0.09$ $16.50\pm 0.10$ $18.30\pm 0.11$
$\mu_{\alpha\star}$ [mas yr-1] 0.044 0.409-0.409 0.120-0.120
$\mu_{\delta}$ [mas yr-1] 0.188-0.188 0.037 0.071
$d$ [kpc] 76 86 76
$v_{\mathrm{los}}$ [km s-1] 290.7-290.7 224.3 247.0-247.0
$\sigma_{\mathrm{los}}$ [km s-1] $9.1\pm 1.2$ $7.9\pm 1.3$ $8.6\pm 0.3$
$N_{\mathrm{tot}}$ 407 440 1046
$N_{\mathrm{DESI}}$ 96 240 1003
$N_{\mathrm{W23}}$ 374 368 416
@@
Table 1: A summary of the properties for Draco, Sextans and UMi. The top eight rows show the central sky coordinates in Equatorial coordinates ($\alpha_{0}$ and $\delta_{0}$), half-light radii ($r_{h}$), systemic proper motions ($\mu_{\alpha}\ast$ and $\mu_{\delta}$), distances ($d$), systemic LOSVs ($v_{\mathrm{los}}$), and intrinsic LOSV dispersions ($\sigma_{\mathrm{los}}$). Distances are from Muñoz et al. (2018). All other data are from Pace et al. (2022). The last three rows provide the numbers of member stars in each dSph. $N_{\mathrm{tot}}$ is the total number of member stars after selections in Section 2.4. $N_{\mathrm{DESI}}$ and $N_{\mathrm{W23}}$ are the numbers of member stars which are observed by DESI and from Walker et al. (2023), respectively. There are overlaps between DESI and Walker et al. (2023). For stars existing in both data set, we use the LOSV and metallicity measurements with smaller errors (see Section 2.3).